# Landsat30-AU: A Vision-Language Dataset for Australian Landsat Imagery

### v2

@@ 5 Benchmark Evaluation
To structure our evaluation, we group the models based on their training domain. The Specialized category comprises remote sensing VLMs, including EarthDial (Soni et al. 2025) and RS-LLaVA (Bazi et al. 2024), and reasoning VLMs, such as GLM-4.1V (GLM-V) (Team et al. 2025c) and MiMo-VL (MiMo) (Team et al. 2025a). The General category consists of foundational models like Qwen2.5-VL (Qwen), Gemma 3 (Gemma3) (Team et al. 2025b), Llama-3.2 (Llama) (Grattafiori et al. 2024), and LLaVA-OneVision (LLaVA) (Li et al. 2024). We ran the two reasoning models in a zero-shot setting, enforcing a maximum output of 8,192 tokens, while the remaining models were evaluated in a one-shot setting.
Furthermore, we fine-tuned two of the general models, Qwen and Llama (creating Qwen-ft and Llama-ft), using LoRA (Hu et al. 2021) on 15% of the respective training data for each task (Landsat30-AU-Cap for captioning and Landsat30-AU-VQA for VQA).
@@ 6 Conclusion
We introduce a new dataset of Landsat satellite optical imagery, comprising data from Landsat 5, 7, 8, and 9. This dataset includes Landsat30-AU-Cap, which provides detailed captions designed to help VLMs align images and text on 30-meter GSD Landsat data for model training and validation. It also features Landsat30-AU-VQA, aimed at assessing VLM capabilities and limitations inherent to Landsat’s 30-meter GSD. Our benchmark evaluation reveals that while fine-tuning is critical for adapting models to this domain, significant challenges remain in complex tasks, highlighting key areas for future VLM development.

### v3

@@ 5 Benchmark Evaluation
To structure our evaluation, we group the models based on their training domain. The Specialized category comprises remote sensing VLMs, including EarthDial (soni2025earthdialturningmultisensoryearth) and RS-LLaVA (rs16091477), and reasoning VLMs, such as GLM-4.1V (GLM-V) (vteam2025glm41vthinkingversatilemultimodalreasoning) and MiMo-VL (MiMo) (coreteam2025mimovltechnicalreport). The General category consists of foundational models like Qwen2.5-VL (Qwen), Gemma 3 (Gemma3) (gemmateam2025gemma3technicalreport), Llama-3.2 (Llama) (grattafiori2024llama3herdmodels), and LLaVA-OneVision (LLaVA) (li2024llavaonevisioneasyvisualtask). We ran the two reasoning models in a zero-shot setting, enforcing a maximum output of 8,192 tokens, while the remaining models were evaluated in a one-shot setting.
Furthermore, we fine-tuned two of the general models, Qwen and Llama (creating Qwen-ft and Llama-ft), using LoRA (DBLP:journals/corr/abs-2106-09685) on 15% of the respective training data for each task (Landsat30-AU-Cap for captioning and Landsat30-AU-VQA for VQA).
@@ 6 Conclusion
We introduce a new dataset derived from optical imagery across the Landsat 5, 7, 8, and 9 missions. This dataset is organized into two subsets: Landsat30-AU-Cap, containing $196,262$ image-captioning pairs, and Landsat30-AU-VQA, comprising $17,725$ human-verified VQA samples. Both components are built from 30-meter resolution Landsat imagery and are curated to facilitate VLM training and validation within this specific domain. Our benchmark evaluation reveals that while fine-tuning is critical for adapting models, significant challenges remain in complex tasks, thereby highlighting key areas for future VLM development.
Acknowledgements
Contribution list:
Sai Ma:
Conceptualization; Data Curation; Methodology; Investigation; Formal Analysis; Writing – Original Draft; Writing – Review & Editing.
Zhuang Li:
Conceptualization; Methodology; Writing – Original Draft; Writing – Review & Editing; Supervision.
John A. Taylor:
Supervision; Conceptualization; Resources; Project Administration.
@@ 52. Read the question and the list of answer options
RQ4: How does fine-tuning affect model robustness across different sensors?
To assess model robustness to sensor-specific variations, we breakdown the VQA test set results by the source Landsat satellite (Landsat 5, 7, 8, and 9). We compare the performance consistency of base models against their fine-tuned counterparts across these four sensors.
As shown in Table 13, base models exhibit significant performance variance across different sensors, indicating a lack of robustness to shifts in sensor characteristics. For instance, the standard Llama model’s accuracy fluctuates from 0.670 on Landsat 8 down to 0.541 on Landsat 5. This inconsistency highlights a critical weakness for real-world applications, where a model must perform reliably on imagery from any available satellite.
Fine-tuning on Landsat30-AU proves crucial for mitigating this issue. After fine-tuning, Llama-ft’s performance stabilizes dramatically, with scores tightly clustered between 0.717 and 0.743 across all four sensors. This alignment reduces the standard deviation in performance across sensors from 0.057 to just 0.010, a testament to improved robustness. Similarly, while the base Qwen model struggles with Landsat 9 imagery (0.679 accuracy), fine-tuning boosts its performance on this sensor to 0.910, bringing it in line with its performance on other sensors. This demonstrates that training on Landsat30-AU, with its diverse, multi-sensor composition, is critical for mitigating sensor-specific biases and developing models that can generalize reliably across the entire 36-year Landsat archive.
[…]
Specialized EarthDial 4B 0.520 0.453 0.514 0.444 0.483 0.041
RS-LLaVA 7B 0.580 0.619 0.491 0.646 0.572 0.068
MiMo 7B 0.761 0.721 0.808 0.707 0.756 0.046
GLM-V 9B 0.684 0.622 0.657 0.559 0.629 0.058
General Qwen 7B 0.732 0.722 0.802 0.679 0.743 0.054
LLaVA 8B 0.588 0.578 0.681 0.543 0.610 0.062
Llama 11B 0.541 0.589 0.670 0.548 0.602 0.059
Gemma 3 12B 0.782 0.743 0.712 0.738 0.736 0.029
General with ft Qwen-ft 7B 0.895 0.919 0.805 0.910 0.871 0.052
Llama-ft 11B 0.717 0.733 0.743 0.720 0.731 0.012
Table 13: VQA performance breakdown by Landsat sensor. The scores represent the accuracy for each model on the subset of the VQA test set corresponding to each satellite. Fine-tuning significantly improves performance consistency across sensors, as indicated by the lower standard deviation ($\sigma$). Bold indicates the best performance in each column.